# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

### v1

@@ Abstract
The paper provides data-dependent bounds on the test error of the Gibbs algorithm in the overparameterized interpolation regime, where low training errors are also obtained for impossible data, such as random labels in classification. The bounds are stable under approximation with Langevin Monte Carlo algorithms. Experiments on the MNIST and CIFAR-10 datasets verify that the bounds yield nontrivial predictions on true labeled data and correctly upper bound the test error for random labels. Our method indicates that generalization in the low-temperature, interpolation regime is already signaled by small training errors in the more classical high temperature regime.
@@ 1 Introduction
Modern learning algorithms can achieve very small training errors on arbitrary data if the underlying hypothesis space is large enough. For reasonable data originating from real-world problems, the chosen hypotheses also tend to have small test errors, a fortunate circumstance, which has given great technological and economic thrust to deep learning. Unfortunately, the same algorithms also achieve very small training errors for data specifically designed to produce very large test errors, such as random labels in classification. Consequently, the hypothesis space and the training error do not suffice to predict the test error. The key to generalization must be more deeply buried in the data.
While not so disquieting to practitioners, this mystery has troubled theoreticians for many years (Zhang et al., 2016; 2021), and it seems safe to say that the underlying mechanisms still have not been completely understood.
@@
We are far from solving this riddle in generality, but for the Gibbs posterior we show how nontrivial bounds on the test error can be recovered from the training data. The Gibbs posterior assigns probabilities, which decrease exponentially with the training error of the hypotheses. The exponential decay parameter $\beta$ can be interpreted as an inverse temperature in an analogy to statistical physics. The Gibbs measure is a sufficient idealization to have tractable theoretical properties, but also the limiting distribution of several concrete stochastic algorithms, here summarized as Langevin Monte Carlo (LMC), including Stochastic Gradient Langevin Dynamics (SGLD), (Gelfand & Mitter, 1991; Welling & Teh, 2011), a popular modern learning algorithm.
@@
When $\beta$ is large and the hypothesis space is rich, these algorithms can reproduce the dilemma described above by achieving very small training errors on data designed to have large test errors. Our paper addresses this interpolation regime of the Gibbs posterior and makes the following three contributions:
@@
  • We give high-probability data-dependent bounds on the test error, both for a hypothesis drawn from the Gibbs posterior and for the posterior mean, assuming that we can freely draw samples from it. These bounds holds for the entire range of temperatures.
  • We weaken the above assumption by showing that the bounds are stable under approximations of the posterior in the total variation and $W_{2}$-Wasserstein metrics. Given sufficient computing resources, this yields bounds for LMC algorithms.
  • Using LMC on real-world classification data (MNIST and CIFAR-10), we obtain non-trivial upper bounds on the test error for true labels and upper bounds for random labels, despite the algorithms achieving very small training error on the random labels.
Our method is based on a combination of the PAC-Bayesian bounds (McAllester, 1999; Alquier, 2021; Rivasplata et al., 2020) with an integral representation of the log-partition function. This makes it possible to bound the logarithm of the density of the Gibbs posterior at a given temperature in terms of empirical averages at higher temperatures. A qualitative conclusion is that generalization in the under-regularized low-temperature regime ($\beta>n$) is already indicated by small training errors in the over-regularized high-temperature regime ($\beta<n$), where $n$ is the number of training examples.
@@ 1.1 Related literature
Many papers address the generalization of the Gibbs algorithm and Langevin Monte Carlo, with special focus on SGLD, which is the most popular algorithm. Most similar to this work is Raginsky et al. (2017), which bounds the distance to the Gibbs posterior and then its generalization error. Their bound, however, applies only to the high temperature regime $\beta<n$.
@@
Several works concentrate on the optimization path of SGLD. Mou et al. (2018) gives both stability and PAC-Bayesian bounds. Pensia et al. (2018) applies the information theoretic generalization bounds of Xu & Raginsky (2017). These ideas are further developed by Negrea et al. (2019), where random subsets of the training data are used to define data-dependent priors. Farghly & Rebeschini (2021) gives time-independent bounds for SGLD, which are further improved by Futami & Fujisawa (2024). Most of the bounds in the above papers are in expectation. The very recent paper of Harel et al. (2025) gives a very elegant argument for Markov chain algorithms based on the second law of thermodynamics. If the invariant distribution is the Gibbs posterior, the bound along the entire optimization path is of order $\sqrt{\beta/n}$ but improvable to $\beta/n$.
@@
Some papers give similar bounds for the Gibbs posterior, roughly of the form $\beta/n$ or $\sqrt{\beta/n}$ (Raginsky et al., 2017; Dziugaite & Roy, 2018; Kuzborskij et al., 2019; Rivasplata et al., 2020) or Maurer (2024) and Harel et al. (2025)). These bounds hold equally for random labels and are therefore vacuous for overparametrized hypothesis spaces in the low temperature regime $\beta>n$. To our knowledge, ours is the only bound for the Gibbs posterior, which is valid in this regime.
@@
Other bounds have been developed for specific algorithms designed to optimize them. The milestone paper by Dziugaite & Roy (2017) is the most prominent example, and (Dziugaite & Roy, 2018) and (Pérez-Ortiz et al., 2021) are also in this category. Our bounds by contrast apply to the Gibbs posterior and LMC in their standard forms.
2 Preliminaries
The relative entropy of two Bernoulli variables with expectations $p$ and $q$ is denoted
We also define the function $\kappa^{-1}:\left[0,1\right]\times\left[0,\infty\right)\rightarrow\left[0,1\right]$ by
@@ 2 Preliminaries
Throughout the following $\left(\mathcal{X},\Sigma\right)$ is a measurable space of data with probability measure $\mu$. The iid random vector $\mathbf{x}\sim\mu^{n}$ is the training sample.
@@
We let $\left(\mathcal{H},\Omega\right)$ be a measurable space of hypotheses, and let $\ell:\mathcal{H\times X}\rightarrow\left[0,\infty\right)$ be a prescribed loss function. Members of $\mathcal{H}$ are denoted $h$ or $g$. We write $L\left(h\right):=\mathbb{E}_{x\sim\mu}\left[\ell\left(h,x\right)\right]$ and $\hat{L}\left(h,\mathbf{x}\right):=\left(1/n\right)\sum_{i}\ell\left(h,x_{i}\right)$ respectively for the true (expected) and empirical error of hypothesis $h\in\mathcal{H}$. The set of probability measures on $\left(\mathcal{H},\Omega\right)$ is denoted $\mathcal{P}\left(\mathcal{H}\right)$.
A stochastic algorithm is a function $\nu:\mathcal{X}^{n}\rightarrow\mathcal{P}\left(\mathcal{H}\right)$, which assigns to a training sample $\mathbf{x}$ a probability measure $\nu\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$. The $KL$-divergence between two probability measures is the function $KL:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\mathbb{E}_{h\sim\rho}\big[\ln\frac{d\rho}{d\nu}\big]$ if $\rho$ absolutely continuous w.r.t. $\nu$, otherwise the value is $\infty$. The total variation distance is defined as $d_{TV}:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\sup_{A\in\Omega}|\rho(A)-\nu(A)|$. The $W_{p}$-Wasserstein distance is $W_{p}(\rho,\nu)=(\inf_{W}\mathbb{E}_{(x,y)\sim W}[\|x-y\|^{p}])^{1/p}$ with the infimum over all probability measures on $\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})$ with $\rho$ and $\nu$ as marginals.
@@
There is an a-priori reference measure $\pi\in\mathcal{P}\left(\mathcal{H}\right)$, called the prior. With a fixed prior, the Gibbs algorithm at inverse temperature $\beta>0$ is the stochastic algorithm $G_{\beta}:\mathbf{x}\in\mathcal{X}^{n}\mapsto G_{\beta}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ defined by
@@
$G_{\beta}\left(\mathbf{x}\right)$ is called the Gibbs-posterior, the normalizing factor
@@
is called the partition function. The motivation for the Gibbs posterior is that it puts larger weights on hypotheses with smaller empirical error.
Then, $\mathbb{E}_{\left(h,\mathbf{x}\right)\sim\rho_{\nu}}\left[\phi\left(h,\mathbf{x}\right)\right]=\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\phi\left(h,\mathbf{x}\right)\right]$ for measurable $\phi:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$. To draw the pair $\left(h,\mathbf{x}\right)$ from $\rho_{\nu}$ we first draw the training sample $\mathbf{x}$, and then sample $h$ from $\nu\left(\mathbf{x}\right)$. The main objective in learning is that the risk $\mathbb{E}_{x\sim\mu}\left[f\left(h,x\right)\right]$ is small with high probability in $\left(h,\mathbf{x}\right)\sim\rho_{\nu}$, where $f$ is some application-dependent loss function, possibly different from $\ell$. In the sequel we will give corresponding guarantees.
3 Bounds for the Gibbs posterior
In this section, we make the idealized assumption that we are free to sample from the Gibbs posterior at any finite $\beta\geq 0$.
3.1 An integral representation of the free energy
@@
Let $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$. Then
$\ln Z_{\beta}(\mathbf{x})=\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]d\gamma\leq\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim G_{\beta_{k-1}}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$

### v2

@@ Abstract
This paper provides data-dependent bounds on the expected error of the Gibbs algorithm in the overparameterized interpolation regime, where low training errors are also obtained for impossible data, such as random labels in classification. The results show that generalization in the low-temperature regime is already signaled by small training errors in the noisier high-temperature regime. The bounds are stable under approximation with Langevin Monte Carlo algorithms. The analysis motivates the design of an algorithm to compute bounds, which on the MNIST and CIFAR-10 datasets yield nontrivial, close predictions on the test error for true labeled data, while maintaining a correct upper bound on the test error for random labels.
@@ 1 Introduction
Modern learning algorithms can achieve very small training errors on arbitrary data if the underlying hypothesis space is large enough. For meaningful data, the chosen hypotheses also tend to have small test errors, a fortunate circumstance, which has given great technological and economic thrust to deep learning. Unfortunately, the same algorithms also achieve very small training errors for data specifically designed to produce very large test errors, such as random labels in classification. In such a situation, which we will loosely refer to as the interpolation regime, the hypothesis space and the training error do not suffice to predict the test error. The key to generalization must be more deeply buried in the data. While not so disquieting to practitioners, this mystery has troubled theoreticians for many years (Zhang et al., 2016, 2021), and it seems safe to say that the underlying mechanisms still have not been completely understood.
@@
We are far from solving this riddle in generality, but for sufficiently close approximations of the Gibbs posterior we show how nontrivial bounds on the test error can be recovered from the training data. The Gibbs posterior assigns probabilities that decrease exponentially with the training error of the hypotheses. The exponential decay parameter $\beta$ can be interpreted as an inverse temperature in an analogy to statistical physics. The Gibbs measure is a sufficient idealization to have tractable theoretical properties, but it is also the limiting distribution of several concrete stochastic algorithms, here summarized as Langevin Monte Carlo (LMC), including Stochastic Gradient Langevin Dynamics (SGLD), (Gelfand and Mitter, 1991; Welling and Teh, 2011), a popular modern learning algorithm.
@@
When $\beta$ is large and the hypothesis space is rich, these algorithms can reproduce the dilemma described above by achieving very small training errors on data designed to have large test errors. Our paper addresses this regime of the Gibbs posterior and makes the following three contributions:
@@
  • We give high-probability data-dependent bounds on the true error, both for a hypothesis drawn from the Gibbs posterior and for the posterior mean, assuming that we can freely draw samples from it. These bounds hold for the entire range of temperatures.
  • We show that these bounds are stable under approximations of the Gibbs posterior in relative entropy. Given enough computing resources, this yields bounds for LMC algorithms, based on known results for non-convex sampling.
  • Existing convergence guarantees for LMC are insufficient for both a practical and rigorous computation of these bounds on real-world problems. A heuristic calibration step, based exclusively on the training data, leads to very close upper bounds on the test error for various neural networks trained with LMC on the MNIST and CIFAR-10 datasets.
The idea underlying our bound is the following. The PAC-Bayesian theorem or its single draw variant (McAllester, 1999; Alquier, 2024; Rivasplata et al., 2020) bounds the generalization error roughly proportional to the logarithm of the posterior density or its posterior expectation (the relative entropy to the prior). The log-density of the Gibbs posterior at inverse temperature $\beta$ has an explicit expression in terms of an integral from 0 to $\beta$ of mean training errors, a fact which seems to have been overlooked in the PAC-Bayesian analysis of generalization (Lemma 3.1 below). Substitution of this integral in the PAC-Bayesian theorem then gives a bound on the generalization error.
As an illustrative example: if the loss $\ell$ has values in $\left[0,1\right]$, and we happen to draw from the Gibbs posterior at $\beta$ a hypothesis $h$ with training error $\hat{L}\left(h,\mathbf{x}\right)=0$, then we have the following bound on the expected error of this hypothesis.
$E_{x}\left[\ell\left(h,x\right)\right]\leq\frac{2\left(A+\ln\left(2\sqrt{n}/\delta\right)\right)}{n},$ (1)
where $A$ is the area in Figure 1, $n$ is the sample size and $\delta$ is the confidence parameter.
Figure 1: The mean training error of the Gibbs posterior is plotted against the inverse temperature. If the training error $\hat{L}\left(h,\mathbf{x}\right)$ is zero, then the log density, $\ln\frac{dG_{\beta}\left(\mathbf{x}\right)}{d\pi}\left(h\right)$, is equal to the area AA.
For the Gibbs posterior this simple, but novel reasoning resolves the dilemma of the interpolation regime. The training error of sufficiently overparametrized systems at a large value of $\beta$ (low temperature) is typically near zero and does not distinguish between easy and hard (e.g. random-label) data, but the mean training losses at small values of $\beta$ (high temperatures) will be quite different, leading to different predictions also at large $\beta$ (low temperature). Paraphrased:
Better generalization in the low-temperature regime is already indicated by smaller training errors in the high-temperature regime.
Figure 2: Mean training error of a fully-connected neural network with 400,000 parameters trained with SGLD on 2,000 examples of MNIST, both with true and random labels.
This principle appears to hold also for the distributions generated by practical LMC algorithms trained on real-world data, as is witnessed by Figure 2. In Section 5.4 we use the area ratio $A/\bar{A}$ and the fact that the true classification error for random labels is $1/2$, to develop a practical method to compute bounds in realistic environments.
After a brief survey of related literature, we review the PAC-Bayesian theorem, introduce the Gibbs posterior and present our bounds, followed by a stability analysis. We conclude with a section describing our experiments. The appendix contains a glossary of notation with section references.
@@ 1.1 Related Literature
Many papers address the generalization of the Gibbs algorithm and Langevin Monte Carlo, with special focus on SGLD, which is the most popular algorithm. Most similar to this work is Raginsky et al. (2017), which bounds the distance of the distribution generated by SGLD to the Gibbs posterior and then the latter’s generalization error. Their bound applies only to the high-temperature regime $\beta<n$, but their convergence guarantees can be combined with our method to give bounds for SGLD in the entire temperature range.
@@
Several works concentrate on the optimization path of SGLD. Mou et al. (2018) give both stability and PAC-Bayesian bounds. Pensia et al. (2018) apply the information-theoretic generalization bounds of Xu and Raginsky (2017). These ideas are further developed by Negrea et al. (2019), where random subsets of the training data are used to define data-dependent priors. Farghly and Rebeschini (2021) give time-independent bounds for SGLD, which are further improved by Futami and Fujisawa (2024). Most of the bounds in the above papers are in expectation. The very recent paper of Harel et al. (2025) gives a very elegant argument for Markov chain algorithms based on the second law of thermodynamics. If the invariant distribution is the Gibbs posterior, the bound along the entire optimization path is of order $\sqrt{\beta/n}$ but improvable to $\beta/n$.
@@
Some papers give similar bounds for the Gibbs posterior, roughly of the form $\beta/n$ or $\sqrt{\beta/n}$ (Raginsky et al., 2017; Dziugaite and Roy, 2018; Kuzborskij et al., 2019; Rivasplata et al., 2020; Maurer, 2024; Harel et al., 2025)). These bounds hold equally for random labels and are therefore vacuous for overparametrized hypothesis spaces in the low temperature regime $\beta>n$. To our knowledge, ours is the only bound for the Gibbs posterior which is valid in this regime.
@@
Other bounds have been developed for specific algorithms designed to optimize them. The milestone paper by Dziugaite and Roy (2017) is the most prominent example; Zhou et al. (2018), Dziugaite and Roy (2018) and Pérez-Ortiz et al. (2021) are also in this category. Our bounds, by contrast, apply to the Gibbs posterior and LMC in their standard forms.
2 The PAC-Bayesian Bound
@@
Throughout the following $\left(\mathcal{X},\Sigma\right)$ is a measurable space of data with probability measure $\mu$. The i.i.d. random vector $\mathbf{x}\sim\mu^{n}$ is the training sample.
@@
$\left(\mathcal{H},\Omega\right)$ is a measurable space of hypotheses, and $\ell:\mathcal{H\times X}\rightarrow\left[0,\infty\right)$ is a prescribed loss function. Members of $\mathcal{H}$ are denoted $h$ or $g$. For a function $f:\mathcal{H}\to\mathbb{R}$, the sup-norm is denoted $f||_{\infty}$. We write $L\left(h\right):=\mathbb{E}_{x\sim\mu}\left[\ell\left(h,x\right)\right]$ and $\hat{L}\left(h,\mathbf{x}\right):=\left(1/n\right)\sum_{i}\ell\left(h,x_{i}\right)$ respectively for the true (expected) and empirical error of hypothesis $h\in\mathcal{H}$. The set of probability measures on $\left(\mathcal{H},\Omega\right)$ is denoted $\mathcal{P}\left(\mathcal{H}\right)$. The relative entropy or Kullback-Leibler-divergence between two probability measures is the function $\text{KL}:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\mathbb{E}_{h\sim\rho}\big[\ln\frac{d\rho}{d\nu}(h)\big]$ if $\rho$ is absolutely continuous w.r.t. $\nu$, otherwise the value is $\infty$. The Rényi-infinity divergence (Rényi, 1961) is $R_{\infty}(\rho,\nu)=\sup_{h\in\mathcal{H}}\ln\frac{d\rho}{d\nu}(h)$ for $\nu,\rho\in\mathcal{P}(\mathcal{H})$. There is an a-priori reference measure $\pi\in\mathcal{P}\left(\mathcal{H}\right)$, called the prior. A stochastic algorithm is a function $\nu:\mathcal{X}^{n}\rightarrow\mathcal{P}\left(\mathcal{H}\right)$, which assigns to a training sample $\mathbf{x}$ a probability measure $\nu\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$.
The following general version of the PAC-Bayesian theorem appears in this form for the first time in Rivasplata et al. (2020). It gives a bound for single hypotheses drawn from the posterior (i) as well as for posterior averages (ii). Appendix B.1 gives an easy proof for the reader’s benefit.
Theorem 2.1.
Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function, and let $\nu$ be a stochastic algorithm such that $\nu\left(\mathbf{x}\right)$ is absolutely continuous w.r.t. $\pi$ for all $\mathbf{x}\in\mathcal{X}^{n}$. Then
(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$
$F\left(h,\mathbf{x}\right)\leq\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(\frac{1}{\delta}\right)$
$\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\text{KL}\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\frac{1}{\delta}$
Here $F$ is a placeholder for a random variable related to the generalization gap, which we want to bound. Suppose $\ell$ has values in $\left[0,1\right]$. With a suitable choice of $F$, we can use (i) above to derive, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right),$ that
$L\left(h\right)$ $\!\!\!\!\!\leq\!\!\!\!\!$ $\hat{L}\left(h,\mathbf{x}\right)+\sqrt{\frac{2\hat{L}\left(h,\mathbf{x}\right)}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\frac{2\sqrt{n}}{\delta}\right)}+\frac{2}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\frac{2\sqrt{n}}{\delta}\right).$ (2)
[…]
3 Bounds for the Gibbs Algorithm
@@
With a fixed prior, the Gibbs algorithm at inverse temperature $\beta>0$ is the stochastic algorithm $G_{\beta}:\mathbf{x}\in\mathcal{X}^{n}\mapsto G_{\beta}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ defined by
@@
$G_{\beta}\left(\mathbf{x}\right)$ is called the Gibbs posterior, the normalizing factor
@@
is called the partition function.
The Gibbs posterior provides a principled, albeit idealized, way to put larger weights on hypotheses with smaller empirical error. It appears as the minimizer of the PAC-Bayesian bound in various forms (see McAllester (1999); Guedj (2019); Alquier (2024)). As $\beta\rightarrow\infty$, the Gibbs posterior concentrates on the set of empirical risk minimizers (Athreya and Hwang, 2010), so the low-temperature regime is particularly interesting.
Evidently $\ln\left(dG_{\beta}\left(\mathbf{x}\right)/d\pi\right)\left(h\right)=-\beta\hat{L}\left(h,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)$. This function has an important integral representation, which is well known from statistical mechanics (see, e.g. Huang (2008)). Despite its simplicity, it seems to have been overlooked in the literature on generalization.
@@
For all $\beta\geq 0$, $\mathbf{x}\in\mathcal{X}^{n}$ and $h\in\mathcal{H}$
$\!\!\!-\ln Z_{\beta}\left(\mathbf{x}\right)$ $=$ $\int_{0}^{\beta}\mathbb{E}_{g\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(g,\mathbf{x}\right)\right]d\gamma$ (3)
$\ln\frac{dG_{\beta}\left(\mathbf{x}\right)}{d\pi}\left(h\right)$ $=$ $\int_{0}^{\beta}\left(\mathbb{E}_{g\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(g,\mathbf{x}\right)\right]-\hat{L}\left(h,\mathbf{x}\right)\right)d\gamma$ (4)
$\text{KL}\left(G_{\beta}\left(\mathbf{x}\right),\pi\right)$ $=$ $\int_{0}^{\beta}\left(\mathbb{E}_{g\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(g,\mathbf{x}\right)\right]-\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\right)d\gamma.$ (5)
Also the function $\beta\mapsto\mathbb{E}_{g\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big]$, if not constant for all $\beta$, is strictly decreasing in $\beta$.